# Clock systems for stochastic and non-deterministic categorical systems theories

### 31 Mar 2026 v1

@@ 1.2 Moore machines for a monad, morphisms of Moore machines, and clocks
An old and central idea of coalgebraic systems (e.g. [8]) is to generalize this by considering a category $\mathcal{S}$ with finite products and a strong monad $M$ on it. Then an $M$-Moore machine $(X,A^{-},A^{+},f^{+},f^{-})$ is given by three objects of $\mathcal{S}$: states $X$, input alphabet $A^{-}$, and output alphabet $A^{+}$, and transition morphisms in $\mathcal{S}$ too:
@@ 2 Systems theories via tangencies
    $TX_{1}$$A_{1}$$TX_{2}$$A_{2}$$\scriptstyle{f_{1}}{\shortmid}$$\scriptstyle{Tyke}$$\scriptstyle{a}$$\scriptstyle{f_{2}}{\shortmid}$
@@ 3 Representable behaviors for discrete-time deterministic Moore machines
$TX_{1}$$A_{1}$$TX_{2}$$A_{2}$$\scriptstyle{f_{1}}{\shortmid}$$\scriptstyle{C}$$\scriptstyle{a}$$\scriptstyle{f_{2}}{\shortmid}$
@@ 6.1 Compositionality of behavior
We lift our behaviors from a functor $\mathsf{Sys}\to\mathsf{Set}$ to a morphism of systems theories from the systems theory of discrete-time stochastic Moore machines to the behavioral (or “Jan Vilma’s-style”, see [12]) systems theory. We do not delve into the double category theory here, instead we simply explain how the behavior functor extends to the rest of the systems theory (interfaces and composition patterns) in concrete terms.
@@
  • A composition pattern $f\colon A\mathrel{\mkern 3.0mu\vbox{\hbox{$\shortmid$}}\mkern-10.0mu{\to}}B$ is sent to the relation (recall that relations are a subclass of spans) between $\mathsf{Chart}((\tilde{\Omega},1),A)$ and $\mathsf{Chart}((\tilde{\Omega},1),B)$ where $a\colon(\tilde{\Omega},1)\to A$ and $b\colon(\tilde{\Omega},1),B$ are related if and only if the following commutes

### 02 Apr 2026 v2

@@ 1.2 Moore machines for a monad, morphisms of Moore machines, and clocks
An old and central idea of coalgebraic systems (e.g. [8]) is to generalize this by considering a category $\mathcal{S}$ with finite products and a strong monad $M$ on it. Then an $M$-Moore machine $(X,A^{+},A^{-},f,f^{\sharp})$ is given by three objects of $\mathcal{S}$: states $X$, input alphabet $A^{-}$, and output alphabet $A^{+}$, and transition morphisms in $\mathcal{S}$ too:
@@ 2 Systems theories via tangencies
    $TX_{1}$$A_{1}$$TX_{2}$$A_{2}$$\scriptstyle{f_{1}}{\shortmid}$$\scriptstyle{Tx}$$\scriptstyle{a}$$\scriptstyle{f_{2}}{\shortmid}$
@@ 3 Representable behaviors for discrete-time deterministic Moore machines
$TX_{1}$$A_{1}$$TX_{2}$$A_{2}$$\scriptstyle{f_{1}}{\shortmid}$$\scriptstyle{Tx}$$\scriptstyle{a}$$\scriptstyle{f_{2}}{\shortmid}$
@@ 6.1 Compositionality of behavior
We lift our behaviors from a functor $\mathsf{Sys}\to\mathsf{Set}$ to a morphism of systems theories from the systems theory of discrete-time stochastic Moore machines to the behavioral (or “Jan Willems-style”, see [12]) systems theory. We do not delve into the double category theory here, instead we simply explain how the behavior functor extends to the rest of the systems theory (interfaces and composition patterns) in concrete terms.
@@
  • A composition pattern $f\colon A\mathrel{\mkern 3.0mu\vbox{\hbox{$\shortmid$}}\mkern-10.0mu{\to}}B$ is sent to the relation (recall that relations are a subclass of spans) between $\mathsf{Chart}((\tilde{\Omega},1),A)$ and $\mathsf{Chart}((\tilde{\Omega},1),B)$ where $a\colon(\tilde{\Omega},1)\to A$ and $b\colon(\tilde{\Omega},1)\to B$ are related if and only if the following commutes